OpenStudy (anonymous):

Use the distributive property to simplify the expression below: -4x-8/-4

OpenStudy (anonymous):

\[\frac{-4x-8}{-4}=\frac{-4(x+2)}{-4}=x+2\]

OpenStudy (anonymous):

-4(x+2)/-4 = (x+2)